Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Bi-Weekly Remedial Activities Update
Report # 24: January 22 – February 9, 2007

« Back to Atlas Tack Superfund Site Fact Sheet

1. ACTIVITIES PERFORMED DURING THE WEEKS OF JAN. 22 – JAN. 26, 2007

  • Completed the creek bed and marsh sampling effort; creek bed excavation  limits generated
  • Completed the excavation activities in the Lagoon, a total volume of approximately 7,956- cy was generated during this excavation activity
  • Backfilled the eastern cell with 2,400-cy of fill material
  • Constructed approximately 500-ft of  haul road through northwest and east fill areas
  • Mobilized materials and equipment south of the barrier: geo-textile, 40-mil liner, bin blocks, sand bags, (2) box culverts, dense grade fill, 8-inch rip-rap, crane mats,  4-inch dry prime pump, trash pumps, 4-inch rigid aluminum piping, 4-inch lay flat hose, steel plates, (2) crawlers, PC-200 excavator, PC-120 excavator and bobcat
  • Completed the construction of the 100-ft X 50-ft transfer station, solidification bins and the mobilization of 20-cy roll-off containers (see attached drawing)
  • Installed  (2) box culverts and berm along the mouth of Boy’s Creek, in the process constructing a reinforced path constructed out of geo-textile, rip-rap and crane mats
  • Installed dedicated sumps north of the barrier and box culvert to control surface water during excavation activities; water pumped from these locations will be re-directed south of the box culvert dam
  • Relocated the crushed brick stockpile located along the southern edge of the Commercial Area to the crushed concrete stockpile
  • Removed additional soils along the northern edge of the northwest fill,  limit was extended horizontally into the Commercial Area by 5-ft
  • Collected confirmatory soil samples from the floor of  the western cell of the Lagoon; awaiting results
  • Begin to transport approximately 1,500-cy of COMM-97 soils to Marion landfill
  • Using a heated pressure washer system, decontaminated all equipment stained by the encountered “blue soils”  

Ongoing Site Activities:

  • Collected waste characterization soil samples from soils originating from the Lagoon Areas for profiling for off-site landfill disposal
    • Last sample collected from the Lagoon is encountered fuels during excavation along the western edge of the Lagoon excavation, product suspected to have originated from the UST
  • Continue water handling activities
  • Continued to inspect trucks prior to exiting the site
  • Continued to provide flaggers to cross trucks across the bike paths along Adams and Tripp Streets
  • Continued stockpile management activities
  • Continue with air monitoring program

2. PROJECTED ACTIVITIES FOR THE WEEK OF FEB. 5 - FEB. 9, 2007

  • Continue to transport approximately 1,500-cy of COMM-97 soils to Marion landfill
  • Receive the first shipment of Enviroblend,  (22) 1-ton sacks and begin treatment of lead and cadmium containing soil
  • Begin to dewater the infiltration and Commercial and Industrial Debris Area (CID) areas, discharging surface water to the town of Fairhaven sewer via manhole at the end of Church St.
  • Continue to excavate marsh soil south of the barrier, solidify and stockpile

Ongoing Site Activities:

  • Continue to collect  pre and post excavation survey data of the excavations
  • Continue with stockpile management activities
  • Continue water handling activities
  • Continued to inspect trucks prior to exiting the site
  • Continued to provide flaggers to cross trucks across the bike paths along Adams and Tripp Streets
  • Continue with air monitoring program

3. AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 11 – 24 January 2007 can be reviewed in the attached table.

Weekly Dust Monitoring Summary
Week #24

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Church Street (µg/m3)
1/11/2007 0 0 10 0
1/12/2007 0 0 10 0
1/15/2007 Precip Precip Precip Precip
1/16/2007 0 0 4 0
1/17/2007 0 46 0 7
1/18/2007 0 0 0 0
1/19/2007 Precip Precip Precip Precip
1/22/2007 NSC NSC NSC NSC
1/23/2007 NSC NSC NSC NSC
1/24/2007 NSC NSC NSC NSC

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*- Readings are for 24hr period

4. EXCAVATION SUMMARY

WEEKS /E 1/19/06 – 1/26/06
TO DATE
ACTIVITIES INITIAL TOTAL QUANTITY INITIAL TOTAL QUANTITY
Commercial Area Excavation Complete 2,655 CY
C.I.D. Excavation None 7,083 CY
East Fill Excavation None 9,046 CY
NW-Fill Area None 8,612 CY
Lagoon Area 1,856 CY 7,956 CY
Total Amount of Soils Excavated to Date: 35,352 - CY
 

Jump to main content.